A 50 kg shell is fired from a cannon on Earth's surface with a potential energy of 196,000 J. What is the shell's maximum height?




1
Expert's answer
2021-02-02T19:46:28-0500

By the definition of the potential energy, we get:


"PE=mgh,""h=\dfrac{PE}{mg}=\dfrac{196000\ J}{50\ kg\cdot9.8\ \dfrac{m}{s^2}}=400\ m."